OpenStudy (thats_kyy):

The incorrect work of a student to solve an equation 2(y + 4) = 4y is shown below: Step 1: 2(y + 4) = 4y Step 2: 2y + 6 = 4y Step 3: 2y = 6 Step 4: y = 3 Which of the following explains how to correct Step 2 and shows the correct value of y? The equation should be y + 4 = 4y after division by 2; y = 5 The equation should be y + 4 = 4y after division by 2; y = 2 2 should be distributed as 2y + 8; y = 4 2 should be distributed as 2y + 8; y = 2

OpenStudy (thats_kyy):

@jim_thompson5910 @UsukiDoll

OpenStudy (usukidoll):

ahahahh the wrong student didn't distribute all the way 2y+8 = 4y 8 = 4y-2y 8 = 2y 4 = y

OpenStudy (thats_kyy):

Your Funny

OpenStudy (usukidoll):

step 1 is an error already... because through distribution we have 2(y+4) = 2y+8 student must have accidentally did 2+4 after doing 2 x y

OpenStudy (thats_kyy):

So The Answer is C

OpenStudy (usukidoll):

ya

OpenStudy (thats_kyy):

Thanks